OpenStudy (anonymous):

Determine if the solution set for the system of equations shown is the empty set, contains one point or is infinite. x + y = 4 x - y = 0

OpenStudy (anonymous):

1. contains one point (x=2, y=2) 2. contains one point (x=2, y= -3) 3.empty set;equations have the same gradient, they are parallel 4. contains one point (x=2 , y=0) 5. infinite solutions; both equations are the same. 6. C (-1,5) 7. C (-2,-3)

OpenStudy (anonymous):

add the two equations together and get 2x=6 and therefore x=3

OpenStudy (imstuck):

No that's not correct, @thomas.charles27. You have to solve these simultaneously to see what the x and the y are.

OpenStudy (imstuck):

x+y=4 x-y=0 For the first one, y = 4-x and for the second one -y=-x+0 or y = x. So sub in x for y in one of the eqautions and you will get x=4-x. Add x to both sides to get 2x=4 so x = 2. Then sub 2 in for x in one of the equations, andd you will get 2+y=4 and y = 2. So the solution set is x=2, y=2.